# TASTE: Text-Aligned Speech Tokenization and Embedding for Spoken Language Modeling

### v2

@@ 4.2 Evaluating Text-Aligned Spoken Language Modeling
4.2.1 Comparing TASLM with Pretrained SLMs
Speech Continuation Evaluation
A typical way to evaluate the pre-trained SLM is by performing conditional generation. Following previous work [11, 22], we use the 3-second prompt speech from the LibriSpeech test-clean. To evaluate the quality of the continuations generated by the SLMs, we employ GPT-4o to assign MOS scores regarding the transcribed speech continuation using ASR, focusing on the semantic coherence of the continuation. In addition, we compute UTMOS as for evaluating the speech quality and naturalness. Last but not least, we conduct human listening test, in which each evaluator is asked to give a MOS score regarding the overall performance of the generated speech continuation. The details of the instructions for GPT-4o and human are in Appendix.
Likelihood-Based Evaluation
Following previous work [11, 33, 22], we also evaluate our SLMs through likelihood-based benchmarks, where the accuracy score is based on whether the model chooses the correct continuation from the two given speech utterances based on its output likelihoods. We adopt two established benchmarks SALMON [25] and spoken StoryCloze [11, 29], which covers the acoustic aspect and the semantic aspect, respectively. Since both benchmarks contain multiple tasks, we report the average accuracy across these tasks within each benchmark for simplicity. The detailed results are in Appendix A.5.1 for the interested readers. We also report the mean of the SALMON and StoryCloze as an overall assessment for both aspects.
Results
The results of TASLM comparing to other pre-trained SLM are in Table 2. First, we highlight that our TASLMs have achieved significantly higher scores on speech continuation across human and machine evaluations; and good performance on the likelihood-based benchmarks. Note that our base language model contains only 1.3 billion parameters, showing the effectiveness of using TASTE for joint modeling. Compared to the cascade method that has the same base model (first row), our $\rm{TASLM}_{\text{emb}}$ scores comparable on GPT-4o but better on human MOS. This indicates that our generated speech is more natural than the cascade one that utilizes TTS for synthesis. Next, our TASLM is the only SLM that not only maintains but even surpasses the performance of its corresponding text-base model. Moreover, we demonstrate that directly using the S3 token for joint modeling following [46] does not yield performance comparable in any aspect. The result further strengthen the intuition behind TASTE (mitigating the length-mismatch during tokenization stage facilitates effective joint spoken language modeling).
Table 3: Evaluation of spoken question answering. Performance across modalities is compared row-wise, where T denotes text and A denotes audio.
Method Mode Web Q. LLaMA-Q.
Mini-Omni $\xrightarrow{}$) T 21.3 39.0
Mini-Omni 0.5B [46] T+A 4.5 11.6
@@ Results
Moshi 7B [6] T+A 26.6 62.3
@@
Llama-Omni-8B [9] T+A 35.5 67.3
@@
TASLM 1B (embed.) T+A 27.1 57.6
We apply few-shot learning to facilitate question answering.
4.2.2 TASLM for Spoken Question Answering
Following [6], we conduct evaluation on spoken question answering to investigate the understanding ability of our TASLM. For this experiment, we use the $\rm{TASLM}_{\text{emb}}$ for simplicity. We compare our pre-trained only SLM with other instruction-finetuned joint SLMs such as Mini-Omni [46], Moshi [6], and Llama-Omni [9]. We use two spoken question answering benchmarks, Web Questions [2] and LLaMA-Questions [30], following [30]. We report the accuracy of answer containment. For fairness, we report not only the performance of the speech-text joint SLMs, but also the base text LLM they used if applicable. Our results indicate that our TASLM is the only method that does not degrade the corresponding text base LLM. We attribute the phenomenon to the effectiveness of our TASTE tokenization for joint speech-text modeling.
5 Conclusion
In this work, we propose Text-Aligned Speech Tokenization and Embedding (TASTE), to facilitate joint speech-text spoken language modeling. By aggregating proper encoder representation through the specialized cross-attention mechanism and taking the ASR model as initialization, we make the speech tokenization text-aligned in an end-to-end manner with no explicit word alignment required. We conduct extensive evaluation on our TASTE tokenizer. Our results show that TASTE allows high quality speech reconstruction at an extremely low bitrate. With our text-aligned speech tokenization and embedding, joint speech-text modeling becomes straightforward and effective. Our experimental results indicate that TASTE enables turning a text LLM into a spoken one with the simple parameter-efficient finetuning technique applied.
Limitation
Several limitations of our current work point to promising avenues for future research. First, neither our TASTE tokenization nor the text-aligned SLM has been optimized for time efficiency; developing a low-latency, streaming variant remains future work. Second, we have evaluated TASTE only on English data—its portability to other languages deserves thorough investigation. Third, although our pretrained SLM generates high-quality continuations, it does not yet support robust turn-taking or instruction-following behavior, both of which are essential for truly interactive systems.
Appendix A Technical Appendices and Supplementary Material
A.1 Related Work
Recent SLMs often require speech tokenization to conduct language modeling with the next prediction objective as the text LLMs. Unlike text, the speech signal is continuous and lengthy, making it difficult to derive proper speech tokenization for spoken language modeling. Common approaches may utilize self-supervised learned (SSL) speech models followed by quantization techniques to extract speech tokens [1, 14, 21, 11, 33]. In addition, audio or speech codec models have also been used for tokenization in recent SLMs [49, 5, 6, 51]. These models are designed for resynthesis, where the speech decoders are jointly learned with the encoders, making them easy to use for spoken language modeling.
With speech tokenization, GSLM [21, 32] first demonstrates the possibility of building an SLM that can generate speech. TWIST [11] further shows that SLM can benefit from initialization with the text-pretrained LLM. With regard to the huge success of text-only LLMs, recent work shifts the focus towards joint speech-text modeling [11, 6, 46]. Challenged by the modality gap between speech and text tokens, different techniques are introduced to facilitate joint modeling. Spirit LM [33] adopts an interleaving strategy; moshi [6] trains its own tokenizer with a reduced token frequency. Moreover, different patterns and strategies such as delayed or sequential generation are introduced for joint modeling, aiming for more reasonable and coherent speech outputs [46].
Despite the increasing demand of joint speech-text modeling [33, 6, 46], we do not find any work discussing the effectiveness of current speech tokenization for it. Moreover, the speech token is often derived with speech or audio-only data  [^2] An exception is CosyVoice [7]. We discuss it in Section 2 since it is related to our method. . Nonetheless, we observe that recent work is trying to mitigate the modality gap by reducing frequency speech token or conducting additional training stage for text-speech alignment. This motivates us to design a speech tokenization that is directly aligned with its text counterpart, tackling the mismatch issue during the tokenization stage.
In the main text, we have mentioned that we utilize a specialized mechanism based on attention to extract and aggregate the encoder representations. We clarify that the text-speech cross-attention mechanism has also been used for fine-grained control of text-to-speech synthesis (TTS). More specifically,  Chen and Rudnicky [3] propose content-style cross-attention to indicate their text-speech cross-attention mechanism that enables style transfer in TTS. Although both utilize specialized text-speech cross-attention mechanism, the design choices and problem formulations are completely different. We attribute of our main novelty to inventing a text-aligned speech tokenization and embedding for joint spoken language modeling, and the text-speech cross attention mechanism is considered and shown to be a clean, effective, and straightforward way of achieving it.
A.2 Tackling the Vocabulary Mismatch
1:Initialization:
4: TASTE embedding $\hat{\bm{z}}=[\hat{z}_{1},\hat{z}_{2},\ldots,\hat{z}_{N}]$
8:     $\bm{v}^{\text{asr}}_{\text{grouped}}\leftarrow[\underbrace{({v}^{\text{asr}}_{1},{v}^{\text{asr}}_{2},{v}^{\text{asr}}_{3})_{1}}_{\text{word}_{1}},\underbrace{({v}^{\text{asr}}_{4})_{2}}_{\text{word}_{2}},\ldots,\underbrace{({v}^{\text{asr}}_{N-1},\bm{v}^{\text{asr}}_{N})_{W}}_{\text{word}_{W}}]$triangleright Group $\bm{v}^{\text{asr}}$ by the words of $\bm{v}$
14:end procedure
15:procedure WordLevelAveraging($\hat{\bm{z}}_{\text{grouped}}$)
16:     $\bar{\bm{z}}\leftarrow[]$triangleright Initialize a new sequence
17:     for  word group index $i\leftarrow 1$ to $W$ do
19:         ${\bar{z}}_{[j:k]}\leftarrow\mathrm{Average}((\hat{z}_{j},\ldots,\hat{z}_{k}))$triangleright Average the word group
20:         append ${\bar{z}}_{[j:k]}$ to $\bar{\bm{z}}$
21:     end for
23:end procedure
25:     $\tilde{\bm{z}}\leftarrow[]$triangleright Initialize a new sequence
26:     for  word group index $i\leftarrow 1$ to $W$ do
28:         $M\leftarrow\mathrm{Length}(({v}^{\text{llm}}_{j},\ldots,{v}^{\text{llm}}_{k}))$triangleright Get the length of the word group.
30:              append $\bar{\bm{z}}[i]$ to $\tilde{\bm{z}}$
31:         end for
32:     end for
33:end procedure
A.3 Training Details
@@
We separate the training process into the two phases: deriving TASTE tokenization and conducting spoken language modeling with TASTE. In the tokenization phase, only the $\mathrm{Aggregator}$, $\mathrm{Quantizer}$, and the $\mathrm{UnitDecoder}$ is trainable. We use the $\mathrm{Adam}$ optimizer and the learning rate is set to 0.0016. The batch size is set to 160 seconds on each of the 8 NVIDIA A6000 GPUs we used. Note that in the first 2 epochs the quantization is not applied. From the beginning of the third epoch, quantization is applied and the $\mathrm{Quantizer}$ starts to be updated. We train the TASTE tokenizer for 5 epochs, which takes about 2 days for learning, with the learning rate gradually decayed.
@@
As for the spoken language modeling training phase, we use the $\mathrm{AdamW}$ optimizer, the $\mathrm{Consine}$ scheduler with the learning rate set to 1e-5. We use 8 Nvidia A6000 GPUs for training. The total batch size summation over the GPUs is set to 768 samples with the gradient accumulation steps set to 2. To reduce the memory overhead and the computational cost, we employ bfloat16 mixed precision during training. Tools such as DeepSpeed [36] and Liger Kernel [13] are also applied to speed up the fine-tuning process.
A.4 Evaluation Details
A.4.1 Human Evaluation
@@
In Table 1, we have shown our result of the MUSRHA human listening test [40]. Following [51], we conduct the evaluation with a hidden reference but without a lowerpass-filtered anchor. We instruct evaluators to rate the perceptual quality of the given samples with respect to the ground truth on a scale of 1 to 100.
@@
Instruction for Human Speech Continuation MOS Evaluation In this test, each sample will contain a short audio clip called "prompt" (3 seconds) and a longer audio clip called "prompt+continuation" ($\sim$15 seconds).
You will be asked to rate the speech quality of the "prompt+continuation" audio clip, specifically focus on the "continuation" part.
The rating should be based on how likely you think that the long audio is a proper continuation of the "prompt" audio.
Specifically, the rating should be based on the following scale:
1: Bad - The "continuation" is not distinguishable or not natural.
2: Poor - The "continuation" is 25% distinguishable.
3: Fair - The "continuation" is 50% distinguishable and natural.
4: Good - The "continuation" is 75% distinguishable and natural.
5: Excellent - The "continuation" is distinguishable, meaningful, and natural.
Distinguishable means that the words in the "continuation" is distinguishable.
Natural means that the "continuation" sounds like a real human voice and a natural continuation of the prompt without considering the content of the speech.
Meaningful means that you can not only distinguish the words but also understand the meaning of the whole "prompt+continuation".
A.4.2 GPT-4o for MOS Evaluation
As introduced in Section 4.2.1, we use GPT-4o to assign MOS scores to the speech continuation results. Here, we describe the detailed procedure. First, whisper-large-v3 is applied to transcribe the generated speech. Then, given the transcription, the text content from the prompt audio, and the instruction template, GPT-4o can produce a score between 1 and 5. The instruction template is provided below:
@@ A.4.2 GPT-4o for MOS Evaluation
A.5 Additional Results
A.5.1 Details on SALMON and StoryCloze
Our detailed results on SALMON and StoryCloze are reported in Table A.5.1. The introductions of the two benchmarks—SALMON and StoryCloze—are described below.
[…]
SALMON for Acoustic Evaluation
StoryCloze for Semantic Evaluation
To evaluate the SLMs’ ability to comprehend semantic coherence and logical reasoning, we employ the spoken version of StoryCloze test (sSC) and the Topic StoryCloze test (tSC) assembled by ([11]). Assessment of narrative understanding involves presenting a four-sentence story setup, followed by two possible endings. These tasks require the model to select the most appropriate conclusion, thereby testing its grasp of causal and temporal relationships within a narrative. Similarly to SALMON, we measure the accuracy of the distinctions based on the likelihood scores.
A.5.2 Report of Standard Deviations
Table 5: Results with standard deviations of Table 1
Method Bitrate Quality Similarity WER $\downarrow$ UTMOS DNSMOS ViSQOL Drtn. Con. Spkr. Sim. MUSHRA Ground Truth 256k 2.1%$\pm$0.07 4.09$\pm$0.32 3.84$\pm$0.26 - - - 76.6$\pm$15.9 Encodec [5] 1500 5.1%$\pm$0.11 1.58$\pm$0.34 3.26$\pm$0.24 3.46$\pm$0.28 0.94$\pm$0.003 0.63$\pm$0.10 - 3000 2.6%$\pm$0.08 2.35$\pm$0.53 3.48$\pm$0.25 3.81$\pm$0.27 0.96$\pm$0.002 0.78$\pm$0.07 25.6$\pm$18.6 500 5.2%$\pm$0.11 1.27$\pm$0.05 2.99$\pm$0.17 2.80$\pm$0.24 0.94$\pm$0.003 0.35$\pm$0.09 - SpeechTokenizer [51] 2000 3.0%$\pm$0.08 3.56$\pm$0.43 3.60$\pm$0.28 3.65$\pm$0.22 0.97$\pm$0.002 0.80$\pm$0.06 53.9$\pm$22.9 4000 2.5%$\pm$0.08 3.90$\pm$0.36 3.76$\pm$0.28 4.03$\pm$0.17 0.98$\pm$0.002 0.92$\pm$0.04 - Mimi [6] 1000 3.1%$\pm$0.09 3.60$\pm$0.37 3.60$\pm$0.30 3.62$\pm$0.26 0.96$\pm$0.002 0.82$\pm$0.06 67.6$\pm$19.8 S3 token (topline) [7] 600 3.0%$\pm$0.09 4.18$\pm$0.27 3.90$\pm$0.24 3.30$\pm$0.26 0.96$\pm$0.002 0.82$\pm$0.09 70.2$\pm$17.0 Text-only (baseline) $\sim$50 5.9%$\pm$0.11 4.31$\pm$0.16 4.11$\pm$0.22 2.44$\pm$0.23 0.57$\pm$0.006 0.78$\pm$0.09 42.6$\pm$27.1 TASTE (ours) $\sim$150 4.4%$\pm$0.11 4.29$\pm$0.18 4.10$\pm$0.22 3.05$\pm$0.26 0.91$\pm$0.003 0.80$\pm$0.08 68.3$\pm$17.1
Table 6: Results with standard deviations of Table 2.
[…]
Table 7: Results with standard deviations of Table 4.2.1.
Method Mode Web Q. LLaMA-Q.
Mini-Omni $\xrightarrow{}$) T 21.3$\pm$0.9 39.0$\pm$2.8
Mini-Omni 0.5B [46] T+A 4.5$\pm$0.5 11.6$\pm$1.8
[…]
Moshi 7B [6] T+A 26.6$\pm$1.0 62.3$\pm$2.8
[…]
Llama-Omni-8B [9] T+A 35.5$\pm$1.1 67.3$\pm$2.7
[…]
TASLM 1B (embed.) T+A 27.1$\pm$1.0 57.6$\pm$2.9
We apply few-shot learning to facilitate question answering.

### v3

@@ A.1.5 Details on SALMON and StoryCloze
SALMON for Acoustic Evaluation
StoryCloze for Semantic Evaluation
To evaluate the SLMs’ ability to comprehend semantic coherence and logical reasoning, we employ the spoken version of StoryCloze test (sSC) and the Topic StoryCloze test (tSC) assembled by Hassid et al. (2023). Assessment of narrative understanding involves presenting a four-sentence story setup, followed by two possible endings. These tasks require the model to select the most appropriate conclusion, thereby testing its grasp of causal and temporal relationships within a narrative. Similarly to SALMON, we measure the accuracy of the distinctions based on the likelihood scores.
Discussion on SALMON and StoryCloze
On SALMON, we observe that our TASLM falls short for background-related attributes (Room, Background) where the samples are added with environmental sounds (echoes, instruments, background noises from FSD50K). Since TASTE tokenization focuses on natural speech and has not being trained on audio with environmental sound and noise, it may not be able to convey such information. However, on the speech related attributes, such as gender and speaker, our TASLM performs much better and is comparable to other SLMs. On StoryCloze, our TASLM successfully retains its semantic capability with effective joint modeling on TASTE, leading to the best results among all pretrained SLMs.
A.1.6 Report of Standard Deviations
Table 10: Results with standard deviations of Table 1
@@ A.1.6 Report of Standard Deviations
Table 11: Results with standard deviations of Table 2.
@@
Table 12: Results with standard deviations of Table 4.
@@
A.2 Training Details
A.2.1 Hyperparameters and Configurations
@@
We separate the training process into the two phases: deriving TASTE tokenization and conducting spoken language modeling with TASTE. In the tokenization phase, only the $\mathrm{Aggregator}$, $\mathrm{Quantizer}$, and the $\mathrm{UnitDecoder}$ is trainable. We use the $\mathrm{Adam}$ (Kingma, 2015) optimizer and the learning rate is set to 0.0016. The batch size is set to 160 seconds on each of the 8 NVIDIA A6000 GPUs we used. Note that in the first 2 epochs the quantization is not applied. From the beginning of the third epoch, quantization is applied and the $\mathrm{Quantizer}$ starts to be updated. We train the TASTE tokenizer for 5 epochs, which takes about 2 days for learning, with the learning rate gradually decayed.
@@
As for the spoken language modeling training phase, we use the $\mathrm{AdamW}$ (Loshchilov and Hutter, 2019) optimizer, the $\mathrm{Consine}$ scheduler with the learning rate set to 1e-5. We use 8 Nvidia A6000 GPUs for training. The total batch size summation over the GPUs is set to 768 samples with the gradient accumulation steps set to 2. To reduce the memory overhead and the computational cost, we employ bfloat16 mixed precision during training. Tools such as DeepSpeed (Rasley et al., 2020) and Liger Kernel (Hsu et al., 2024) are also applied to speed up the fine-tuning process of the SLM.
A.3 Evaluation Details
A.3.1 Human Evaluation
@@
In Table 1, we have shown our result of the MUSRHA human listening test (Series, 2014). Following Zhang et al. (2024), we conduct the evaluation with a hidden reference but without a lowerpass-filtered anchor. We instruct evaluators to rate the perceptual quality of the given samples with respect to the ground truth on a scale of 1 to 100.
@@
Instruction for Human Speech Continuation MOS Evaluation In this test, each sample will contain a short audio clip called "prompt" (3 seconds) and a longer audio clip called "prompt+continuation" ($\sim$15 seconds). You will be asked to rate the speech quality of the "prompt+continuation" audio clip, specifically focus on the "continuation" part. The rating should be based on how likely you think that the long audio is a proper continuation of the "prompt" audio. Specifically, the rating should be based on the following scale: 1: Bad - The "continuation" is not distinguishable or not natural. 2: Poor - The "continuation" is 25% distinguishable. 3: Fair - The "continuation" is 50% distinguishable and natural. 4: Good - The "continuation" is 75% distinguishable and natural. 5: Excellent - The "continuation" is distinguishable, meaningful, and natural. Distinguishable means that the words in the "continuation" is distinguishable. Natural means that the "continuation" sounds like a real human voice and a natural continuation of the prompt without considering the content of the speech. Meaningful means that you can not only distinguish the words but also understand the meaning of the whole "prompt+continuation".
A.3.2 GPT-4o for MOS Evaluation
As introduced in Section 5.1.2, we use GPT-4o to assign MOS scores to the speech continuation results (Chiang and Lee, 2023; Lin et al., 2024). Here, we describe the detailed procedure. First, whisper-large-v3 is applied to transcribe the generated speech. Then, given the transcription, the text content from the prompt audio, and the instruction template, GPT-4o can produce a score between 1 and 5. The instruction template is provided below:
@@ A.3.2 GPT-4o for MOS Evaluation
A.4 Tackling the Vocabulary Mismatch
1:Initialization:
4: TASTE embedding $\hat{\bm{z}}=[\hat{z}_{1},\hat{z}_{2},\ldots,\hat{z}_{N}]$
8:  $\bm{v}^{\text{asr}}_{\text{grouped}}\leftarrow[\underbrace{({v}^{\text{asr}}_{1},{v}^{\text{asr}}_{2},{v}^{\text{asr}}_{3})_{1}}_{\text{word}_{1}},\underbrace{({v}^{\text{asr}}_{4})_{2}}_{\text{word}_{2}},\ldots,\underbrace{({v}^{\text{asr}}_{N-1},\bm{v}^{\text{asr}}_{N})_{W}}_{\text{word}_{W}}]$ $\triangleright$ Group $\bm{v}^{\text{asr}}$ by the words of $\bm{v}$
14:end procedure
15:procedure WordLevelAveraging($\hat{\bm{z}}_{\text{grouped}}$)
16:  $\bar{\bm{z}}\leftarrow[]$ $\triangleright$ Initialize a new sequence
17:  for word group index $i\leftarrow 1$ to $W$ do
19:   ${\bar{z}}_{[j:k]}\leftarrow\mathrm{Average}((\hat{z}_{j},\ldots,\hat{z}_{k}))$ $\triangleright$ Average the word group
20:   append ${\bar{z}}_{[j:k]}$ to $\bar{\bm{z}}$
21:  end for
23:end procedure
25:  $\tilde{\bm{z}}\leftarrow$[] $\triangleright$ Initialize a new sequence
26:  for word group index $i\leftarrow 1$ to $W$ do
28:   $M\leftarrow\mathrm{Length}(({v}^{\text{llm}}_{j},\ldots,{v}^{\text{llm}}_{k}))$ $\triangleright$ Get the length of the word group.
30:     append $\bar{\bm{z}}[i]$ to $\tilde{\bm{z}}$
31:   end for
32:  end for
33:end procedure
A.5 The Weighted Sum Mechanism for Modality Fusion
The speech decoder takes the text-aligned speech embedding $\hat{\bm{z}}$ and the text embedding $\bm{v}$ as input conditions. As depicted in Figure 2, in practice we perform a weighted sum mechanism to obtain the final fused embedding $\bm{z}_{\text{joint}}$. The weights to fuse the two modalities are represented by two learnable parameters, denoted as $w_{\text{sp}}$ and $w_{\text{txt}}$. The whole process of modality fusion can be denoted as follows:
$\hat{\bm{z}}^{\prime},\bm{v}^{\prime}=\mathrm{normalize}(\hat{\bm{z}})\text{,}\mathrm{normalize}(\bm{v})\text{,}[p_{\text{sp}},p_{\text{txt}}]=\mathrm{softmax}([w_{\text{sp}},w_{\text{txt}}]),z_{\text{joint}}=p_{\text{sp}}\cdot\hat{\bm{z}}^{\prime}+p_{\text{txt}}\cdot\bm{v}^{\prime}\text{.}$
A.6 Discussion on the Usage of LLM
We discuss our usage of LLM following the conference’s policy. We use an AI assistant (ChatGPT specifically) to polish English prose, including grammar correction, wording refinements, consistent terminology and hyphenation, and minor restructuring for clarity and flow. The assistant suggests alternative phrasings, section bridges, and standard disclosure/impact wording based on author-provided content. It does not generate novel ideas, claims, analyses, figures, code, or results, and it does not access proprietary data. All technical content and conclusions are our own, and we review and edit all AI-assisted text and take full responsibility for the final manuscript.